# Approximation and FPT Algorithms for Finding DM-Irreducible Spanning Subgraphs

### v2

@@ 4.1 FPT Algorithm for Unweighted SCSS
The algorithm utilizes another well-known characterization of strongly connected digraphs. For a digraph $G=(V,E)$, an ear decomposition of $G$ is a sequence $\mathcal{E}=(P_{0},P_{1},\dots,P_{f})$ of edge-disjoint paths and cycles in $G$ with $E=\bigcup_{i=0}^{f}E(P_{i})$ such that
@@
  • $P_{i}$ is either a path between two different vertices in $\bigcup_{j=0}^{i-1}V(P_{j})$ whose inner vertices are in $V\setminus\bigcup_{j=0}^{i-1}V(P_{j})$, or a cycle intersecting exactly one vertex in $\bigcup_{j=0}^{i-1}V(P_{j})$.
Each path or cycle $P_{i}$ is called an ear. An ear is called long if it is of length at least $3$.
Lemma 4.1 (cf. [11, Theorem 2.2.1]).
@@
The idea of the FPT algorithm for UnweightedSCSS in [2] is sketched as follows. It first computes an ear decomposition $\mathcal{E}=(P_{0},P_{1},\dots,P_{f})$ of the input graph $G$ so that the set of long ears is inclusion-wise maximal (i.e., so that no more long ear can be taken in addition to them, but the number of long ears in $\mathcal{E}$ is not necessarily maximized). If the number of long ears is large enough (depending on the parameter $k$), then the ear decomposition $\mathcal{E}$ gives a strongly connected spanning subgraph that is small enough by removing all the ears of length $1$, and the answer is yes. Otherwise, let $X$ be the set of vertices intersecting a long ear in $\mathcal{E}$. Under the assumption that $\mathcal{E}$ does not immediately give a solution, we can show that $|X|=O(k)$. Also, by the maximality of long ears in $\mathcal{E}$, the vertices in $V\setminus X$ are added by ears of length $2$, and this is basically true for some minimum strongly connected spanning subgraph. Thus, we can test the existence of a solution by an exhaustive search on the subgraphs of $G$ on the vertex set $X$ (plus a small subset), where we have only $2^{O(k^{2})}$ candidates.
@@ 4.2 FPT Algorithm for Unweighted DMISS
First, we give an alternative characterization of DM-irreducible balanced bipartite graphs by ear decompositions. An ear decomposition of a bipartite graph $G=(V^{+},V^{-};E)$ is defined in the same way as above, and we call it odd proper if
@@
Lemma 4.2 (cf. [11, Theorem 2.4.17] or [23, Theorem 4.1.6]).
@@
An ear is called trivial if it consists of a single edge, and nontrivial otherwise. Also, an ear is called long if it is of length at least $5$. [^3] This redefinition is reasonable by considering the relation between the strong connectivity of a digraph and the DM-irreducibility of a balanced bipartite graph (cf. [3, Section 2.3])
@@
  • $\{P_{1},P_{2},\dots,P_{s}\}$ is the set of long ears in $\mathcal{E}$, which is inclusion-wise maximal (i.e., no more long ear can be taken in addition to them, where $s$ is not necessarily maximized), and

### v3

@@ 4.1 FPT Algorithm for Unweighted SCSS
The FPT algorithm utilizes another well-known characterization of strongly connected digraphs. For a digraph $G=(V,E)$, an ear decomposition of $G$ is a sequence $\mathcal{E}=(P_{0},P_{1},\dots,P_{f})$ of edge-disjoint paths and cycles in $G$ with $E=\bigcup_{i=0}^{f}E(P_{i})$ such that
@@
  • $P_{i}$ $(i=1,2,\dots,f)$ is either a path between two different vertices in $\bigcup_{j=0}^{i-1}V(P_{j})$ whose inner vertices are in $V\setminus\bigcup_{j=0}^{i-1}V(P_{j})$, or a cycle intersecting exactly one vertex in $\bigcup_{j=0}^{i-1}V(P_{j})$.
Each path or cycle $P_{i}$ is called an ear. An ear is called long if it is of length at least $3$, where the length is evaluated by the number of edges.
Lemma 4.1 (cf. [12, Theorem 2.2.1]).
@@
The idea of the FPT algorithm for UnweightedSCSS is sketched as follows. It first computes an ear decomposition $\mathcal{E}=(P_{0},P_{1},\dots,P_{f})$ of the input graph $G$ so that the set of long ears is inclusion-wise maximal (i.e., the set of long ears in $\mathcal{E}$ is not a proper subset of that in any ear decomposition $\mathcal{E}^{\prime}$ of $G$, but the number of long ears in $\mathcal{E}$ is not necessarily maximized). If the number of long ears is large enough (depending on the parameter $k$), then the ear decomposition $\mathcal{E}$ gives a strongly connected spanning subgraph that is small enough by removing all the ears of length $1$, and the answer is yes. Otherwise, let $X$ be the set of vertices intersecting a long ear in $\mathcal{E}$. Under the assumption that $\mathcal{E}$ does not immediately give a solution, we can show that $|X|=O(k)$. Also, by the inclusion-wise maximality of long ears in $\mathcal{E}$, the vertices in $V\setminus X$ are added by ears of length $2$, and this is basically true for some minimum strongly connected spanning subgraph. Thus, we can test the existence of a solution by an exhaustive search on the subgraphs of $G$ on the vertex set $X$ (plus a small subset), where we have only $2^{O(k^{2})}$ candidates.
@@ 4.2 FPT Algorithm for Unweighted DMISS
First, we give an alternative characterization of DM-irreducible balanced bipartite graphs by ear decompositions. An ear decomposition of a bipartite graph $G=(V^{+},V^{-};E)$ is defined in the same way, and we call it odd proper if
@@
Lemma 4.2 (cf. [12, Theorem 2.4.17] or [24, Theorem 4.1.6]).
@@
An ear is called trivial if it consists of a single edge, and nontrivial otherwise. Also, an ear is called long if it is of length at least $5$. [^3] This redefinition is reasonable by considering the relation between the strong connectivity of a digraph and the DM-irreducibility of a balanced bipartite graph (cf. [3, Section 2.3]); a path of length $\ell$ in the former corresponds to a path of length $12\ell-$ 1 in the latter (as a vertex in the former is split into two adjacent vertices in the latter).
@@
  • $\{P_{1},P_{2},\dots,P_{s}\}$ is the set of long ears in $\mathcal{E}$, which is inclusion-wise maximal (i.e., the set of long ears in $\mathcal{E}$ is not a proper subset of that in any ear decomposition $\mathcal{E}^{\prime}$ of $G$, but $s$ is not necessarily maximized), and